OpenStudy (anonymous):

what is the answer to 5-3(4+x)-4x

OpenStudy (anonymous):

Are you trying to simplify this expression?

OpenStudy (anonymous):

5-12-3x-4x=0 -7-7x=0 -7=7x

OpenStudy (anonymous):

x = -1

OpenStudy (anonymous):

Once you get -7=7x, you simply divide to isolate the variable

OpenStudy (anonymous):

@gallaxhar is correct

OpenStudy (anonymous):

It's x = -1. I see @gallaxhar has got it. So, I agree with gallaxhar. (:

OpenStudy (anonymous):

Solve for x: -7 (1+x) = 0 Divide both sides by a constant to simplify the equation. Divide both sides of -7 (1+x) = 0 by -7: 1+x = 0/(-7) Simplify 0/(-7). 0/(-7) = 0: 1+x = 0 Isolate terms with x to the left hand side. Subtract 1 from both sides: x+(1-1) = -1 Look for two terms that sum to zero. 1-1 = 0: Answer: x = -1

OpenStudy (anonymous):

True that.